OpenStudy (anonymous):

Write the quadratic equation in general form and then choose the value of "b." (2x - 1)(x + 5) = 0

myininaya (myininaya):

So you need to do some multiplying

myininaya (myininaya):

\[(ux+v)(ex+d)=ux(ex+d)+v(ex+d)=ux(ex)+ux(d)+v(ex)+v(d)\] \[uex^2+udx+evx+vd\] Combine like terms: \[uex^2+(ud+ev)x+vd \]

OpenStudy (anonymous):

Is it simpler ??

OpenStudy (anonymous):

9?

myininaya (myininaya):

General form is \[ax^2+bx+c=0\] That is why you need to multiply

OpenStudy (anonymous):

@zse can you multiply them ??

OpenStudy (anonymous):

i got 9

OpenStudy (anonymous):

Yes you are right..